OpenStudy (anonymous):

Solve the system using elimination. 4x + 2y = –10 4x – 2y = –14

OpenStudy (anonymous):

Ok, lets get rid of the y first and solve for x 4x+2y=-10 4x-2y=-14 <--- You add the like terms 8x=-24 x=-3 Now you plug x= -3 back into either equation and solve for y.

OpenStudy (anonymous):

subtract the equations 4x+2y=-10 4x-2y=-14 8x=24 x=3

OpenStudy (anonymous):

A.(–3, 1) B.(–2, –1) C.(2, 4) D.(1, –3)

OpenStudy (anonymous):

A is the answer.